OpenStudy (anonymous):

What is the answer to (4)^2-(1/2)^2 ?? thank you

OpenStudy (anonymous):

Well we know that \[ x^2=x*x\] So lets apply it to this case \[4^2-\left(\frac{1}{2}\right ) ^2=4*4-\frac{1}{2}*\frac{1}{2}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Btw this aint the full answer. You still gotta solve it